Part A In a process cost system, units to be accounted for in a department are...

Part A

In a process cost system, units to be accounted for in a department are equal to the

units in the beginning inventory plus the units started or transferred into the department.
number of units transferred out of the department.
ending inventory plus the units started or transferred into the department.

number of units started or transferred into the department.

Part B

Sheffield and Miller Manufacturing is trying to determine the equivalent units for conversion costs with 11900 units of ending work in process at 80% completion and 32300 physical units. There are no beginning units in the department. Conversion costs occur evenly throughout the entire production period. What are the equivalent units for conversion costs for the current period?

44200.
40375.
29920.

9520.
Answer #1

Part A

The answer is units in the beginning inventory plus the units started or transferred into the department.

Part B

Units transferred out = 32,300 - 11,900 = 20,400

Equivalent units for ending inventory = 11,900 * 80% = 9,520

Equivalent units for conversion costs for the current period = 20,400 + 9,520

= 29,920